# Efficient Uncertainty Propagation in Bayesian Two-Step Procedures

### v1

@@ 3.1.2 Results of the Simulation Study
Figure 4 displays the number of log-probability and gradient log-probability evaluations required to approximate the posteriors of all $m=100$ imputed datasets ($N=100$, $p=10$), where in the original dataset 15% of rows contained missing values. The results clearly demonstrate that our proposed method substantially reduces the number of required evaluations compared to the brute-force MCMC baseline – both in terms of log-probability and gradient computations. The greatest efficiency is achieved when combining PSIS and IWMM (PSIS+IWMM), resulting in the lowest computational cost while preserving inference quality and adequately propagating uncertainty across the imputed datasets.
Figure 5 presents the number of MCMC runs required to approximate the posteriors of all $m=100$ imputed datasets. The number of MCMC runs corresponds to the number of distinct proposal distributions needed for successful approximation of all target posterior distributions using our method, compared with the brute-force approach of running MCMC for all imputed datasets. The left panel of the figure shows results for the lower-dimensional case with $p=10$ predictors. Here, depending on the proportion of missing rows in the original dataset, between 1 and 55 MCMC runs are sufficient – highlighting the efficiency gains relative to running MCMC for all 100 datasets individually. As expected, the number of required MCMC runs increases with the proportion of missing data. Higher rates of missingness lead to greater variation among the imputed datasets and thus more diverse posterior distributions, making approximation more difficult for PSIS:single, PSIS:mixture and PSIS+IWMM. In the high-dimensional case with $p=30$ predictors (right panel), we observe that PSIS alone – whether based on a single proposal or a mixture – fails when the proportion of rows containing missing values is large. Nonetheless, PSIS+IWMM continues to offer substantial efficiency gains in most settings using standard priors. However, under the more challenging regularized horseshoe prior and 30% missing data, PSIS+IWMM also struggles to reduce the number of required MCMC runs effectively.
@@
Figure 5: Number of MCMC runs required to approximate the posteriors of $m=100$ imputed datasets with $N=100$ observations, across varying model complexities ($p=10$ or $p=30$, columns) and prior specifications (standard brms priors vs. regularized horseshoe, rows). Representative datasets were selected using the Medoids procedure. The dashed horizontal reference line at 100 indicates comparable performance to the brute-force MCMC approach.
While the previous figure focused on the number of MCMC runs, Figure 6 reports the ratio of log-probability evaluations needed by the iterative algorithm relative to the number required with brute-force MCMC (baseline). A ratio close to 1, indicated by the horizontal reference line, implies that the computational effort in terms of log-probability evaluations is comparable to performing 100 separate MCMC runs. For the lower-dimensional setting (left column), the iterative algorithm requires fewer than half as many evaluations, demonstrating substantial computational savings. Results are similar for both the standard priors and the regularized horseshoe prior, though PSIS+IWMM performs slightly better with the standard priors. In contrast, for the higher-dimensional case (right column), PSIS+IWMM yields efficiency gains only when standard priors are used. In the most challenging scenario – 30% missing data and the horseshoe prior – the number of log-probability evaluations exceeds that of the brute-force MCMC baseline, indicating that the method becomes less efficient in this setting.
Figure 6: Ratio of number of evaluations of the log-probability (on a log-scale) compared to running MCMC separately necessary to approximate the posteriors of $m=100$ imputed datasets with size $N=100$ and $p$ observations (in the columns) using either the brms standard priors or the regularized horseshoe prior (in the rows). The representative datasets are selected using the Medoids procedure. The dashed horizontal reference line at 1 indicates comparable performance to the brute-force MCMC approach.
A similar pattern is observed when examining the ratio of gradient log-probability evaluations, as shown in Figure 7. Unlike log-probability evaluations, gradient evaluations are only required during HMC sampling and are not needed for the computation of importance ratios in PSIS or IWMM. Consequently, for the inference methods based on our iterative algorithm (PSIS:single, PSIS:mixture, and PSIS+IWMM), the gradient evaluations originate exclusively from the HMC runs used to generate proposal distributions. This structural difference explains why these methods inherently require fewer or, at most, the same number of gradient evaluations as the brute-force MCMC approach. For the standard priors (top row in Figure 7), PSIS+IWMM shows even greater improvements in terms of gradient efficiency compared to the gains observed for log-probability evaluations. In the more challenging setting with high dimensionality and the horseshoe prior (bottom right), the performance of PSIS+IWMM is comparable to that of PSIS:single and PSIS:mixture, indicating that in this case, the added complexity of PSIS+IWMM does not lead to substantial gains.
@@
Figure 7: Ratio of number of evaluations of the gradient log-probability (on a log-scale) compared to running MCMC separately necessary to approximate the posteriors of $m=100$ imputed datasets with size $N=100$ and $p$ observations (in the columns) using either the brms standard priors or the regularized horseshoe prior (in the rows). The representative datasets are selected using the Medoids procedure. The dashed horizontal reference line at 1 indicates comparable performance to the brute-force MCMC approach.
@@
Overall, we observe that our proposed method offers substantial improvements in terms of number of log-probability and gradient evaluations – an indicator of computational cost – across most settings. For simpler priors, which result in relatively straightforward posterior distributions, the iterative approach combining both PSIS and IWMM significantly outperforms the baseline MCMC approach. Even in cases involving more complex posterior shapes, such as those induced by the regularized horseshoe prior, the proposed method continues to outperform alternative inference strategies. The only exception arises in high-dimensional settings (i.e., $p=30$ covariates) combined with a high proportion of missing data (i.e., 30% of rows with missing values), where the method does not yield improvements over the brute-force MCMC approach.
@@ 3.2 Multiple Imputation – Case Study
To evaluate the proposed method on real-world data, we examine a travel time prediction problem. Our case study is inspired by the work of Huang et al. (2020) and utilizes a dataset provided by the NYC Taxi and Limousine Commission (TLC) [62], containing yellow taxi trip records in New York City for the year 2016. The dataset includes information such as pickup and dropoff latitude and longitude, timestamps, number of passengers, and vendor ID. Additionally, it contains variables that are only known after trip completion, such as trip distance, trip duration, and payment amount.
We focus on a short-term travel time prediction task, in line with the forecasting problem of Huang et al. (2020). For this purpose, we use only variables available prior to the start of the trip: pickup and dropoff latitude and longitude. In addition, we employ the Open Source Routing Machine (OSM) [63], which uses OpenStreetMap data for New York City to estimate trip duration and distance based on the trip’s start and end locations. These six variables serve as predictors in a Bayesian linear regression model, with the trip duration as the target variable. The model includes a log transformation of the target and scaling of all predictors, and we use the default priors provided by the brms [52] package.
For our analysis, we select data from a single day – Monday, March 7, 2016. The dataset is divided into 24 subsets, one for each hour of the day, based on the trip start time. Depending on the hour, the number of observations ranges from approximately 2,000 to 24,000. Each hourly subset contains some missing values, although the target variable (trip duration) is always complete. The missing data in the predictors are partly due to failures in the OSM estimation process. The proportion of missing values per hourly dataset ranges from 1.1% to 1.7%.
To account for uncertainty due to missing data, we perform multiple imputation using MICE [50], generating $m=20$ imputed datasets for each hour. We then apply the iterative algorithm described in Section 2.3 to approximate the marginalized posterior distributions of the model parameters, using both our full approach (PSIS+IWMM) and a simplified variant without the IWMM step (PSIS:single) that employs a single proposal distribution. As demonstrated in the simulation study in Section 3.1, incorporating a mixture proposal in the multiple imputation context did not offer significant advantages over using a single proposal. Therefore, we adopted the simpler approach in this analysis. For comparison, we also perform HMC sampling (brute-force MCMC) using brms [52, 53] on each of the 20 imputed datasets individually.
Figure 8(b) presents the marginalized posterior distributions for one of these hourly datasets, specifically from 12–1 pm. The red line represents the density obtained from the brute-force MCMC approach, marginalized over all $m=20$ imputed datasets, while the blue line shows the corresponding marginalized posterior density from our iterative method using PSIS and IWMM. The close alignment in shape and location of the distributions indicates that our method closely approximates the gold-standard HMC results.
@@ 3.3 Surrogate Models – Simulation Study
To evaluate the performance of the iterative algorithm in a different context, we conduct a case study based on surrogate modeling. This study replicates and extends the work of Reiser et al. (2025). There, a two-step procedure for uncertainty quantification in surrogate models was developed and tested in multiple experiments. One of these experiments is now re-examined here using the newly developed approach.
@@ 3.3.1 Simulation Setup
~$\widetilde{\mathcal{M}}_{\rm logistic}(\theta,\tau)=\frac{\tau_{1}}{1+\exp(-\tau_{2}(\theta-\tau_{3}))}+\tau_{4}\,,$ (31)
@@
For the training phase of the surrogate, we assign normal priors to the surrogate parameters with means equal to the true values $(2,10,0,-1)$ and the standard deviations set to $(1,10,1,1)$.
@@
~$\widetilde{\mathcal{M}}_{\rm PCE}(\theta,\tau)=\sum_{i=0}^{d}\tau_{i}\psi_{i}(\theta)\,,$ (32)
where the trainable parameters are $\tau=[\tau_{0},\tau_{1},\dots,\tau_{d}]$, and $\psi_{i}$ denote the Legendre polynomials [67]. For this study, we choose a polynomial degree of $d=5$. Each coefficient $\tau_{i}$ is assigned a normal prior, $\tau_{i}\sim\mathcal{N}(0,5)$.
Surrogate Training Step In the initial step, we generate $N_{T}=10$ training data points using the simulator. The input values $\{\theta_{T}^{(n)}\}_{n=1}^{N_{T}}$ are selected as $N_{T}$ equally spaced points in the interval $[-1,1]$. The corresponding outputs $\{y_{T}^{(n)}\}_{n=1}^{N_{T}}$ are sampled from a normal distribution centered at the true simulator values with a standard deviation $\sigma=0.01$, i.e.,
@@
To infer the surrogate parameters, we use MCMC implemented via the probabilistic programming language Stan [53], using the cmdstanr package in R [68]. Two chains are run for 1050 iterations each (including 1000 warm-up and 50 sampling iterations), resulting in $m=100$ posterior draws $\{\tau^{(i)}\}_{i=1}^{m}\sim p(\tau\mid D_{T})$. The relatively small number of posterior draws is chosen to facilitate a more direct comparison with the results from the missing data simulation study in Section 3.1.
@@ 3.3.2 Results of the Simulation Study
We first examine the number of MCMC runs required by the iterative algorithm, which are used to obtain the posterior distributions serving as proposal distributions in PSIS and IWMM. Since no substantial differences between selection strategies can be observed, presented results are from the log-lik method only. Additional results for other selection methods can be found in Appendix D. Figure 9 displays the total number of MCMC runs, categorized by inference method and surrogate type (logistic and PCE). For the logistic surrogate, we observe that PSIS:single requires the highest number of MCMC runs, with a median of 37. In contrast, for PSIS:mixture only about 19 posterior distributions need to be fitted using MCMC. In most cases, the PSIS+IWMM method finishes within 1 or 2 posteriors fitted with MCMC, effectively approximating all posterior distributions and thus requiring a significantly smaller number of MCMC runs. The PCE surrogate represents a more challenging scenario, leading to greater variability across runs. Both methods PSIS:single and PSIS:mixture perform similarly, with a median number of MCMC runs of 70 and 63 for the single and mixture proposals, respectively. Despite some outlier cases, the combined the PSIS+IWMM inference method consistently outperforms the other two in terms of minimizing the number of MCMC runs with a median of 5.
@@
For the logistic surrogate, the PSIS:single requires slightly more log-probability evaluations than the baseline MCMC, but significantly fewer gradient evaluations, which are typically more computationally expensive. The PSIS:mixture inference method, as well as PSIS+IWMM, succeed in approximating all posterior distributions with fewer evaluations of both log-probability and gradient evaluations. In contrast, the PCE surrogate poses a greater challenge, as the posterior distributions differ more substantially. In the worst case – observable in PSIS:single – the iterative algorithm requires more log-probability evaluations than baseline MCMC. This occurs because the algorithm needs many iterations to evaluate all posterior distributions. During each iteration, importance ratios are computed for each of the remaining posteriors. If PSIS fails in many of these cases, a significant number of importance ratio calculations are only used to determine that PSIS does not work for this particular proposal-target combination, rather than approximating the target distribution. As a result, the number of log-probability evaluations exceeds the ones of running MCMC separately. In contrast, the number of gradient evaluations required by the new method cannot exceed those needed for running brute-force MCMC, as they are only necessary for calculating the proposal in the iterative algorithm with HMC, and there can be at most $m=100$ different proposals. The PSIS:mixture method substantially improves PSIS performance compared to the single proposal with only having a few more log-probability evaluations than the MCMC method. The combined PSIS+IWMM approach again results in the fewest gradient evaluations, indicating that only a small number of posteriors needed to be computed using HMC. The high number of log-probability evaluations in this method arises from applying the IWMM algorithm (see Section 2.2) in each iteration of the iterative algorithm for multiple different target distributions.
@@
In conclusion, the proposed method enhances the efficiency of posterior approximation for computing the marginalized posterior via the Monte Carlo estimate. In the simpler setting – using the logistic surrogate, which incorporates the true simulator – the method can approximate all posterior distributions with only one or two expensive HMC runs. In the more complex scenario involving the PCE surrogate, the method faces greater challenges but still achieves notable reductions in computational effort. The number of gradient log-probability evaluations can be reduced to less than 14% of those required by the brute-force MCMC approach, as shown in Table 1. Since gradient evaluations typically dominate the overall computational cost, our proposed method achieves improved efficiency even in more complex settings.

### v2

@@ 3.1.2 Results of the Simulation Study
Figure 5 displays the number of log-probability and gradient log-probability evaluations required to approximate the posteriors of all $m=100$ imputed datasets ($N=100$, $p=10$), where in the original dataset 15% of rows contained missing values. The results clearly demonstrate that our proposed method substantially reduces the number of required evaluations compared to the brute-force MCMC baseline — both in terms of log-probability and gradient computations. The greatest efficiency is achieved when combining PSIS and IWMM (PSIS+IWMM), resulting in the lowest computational cost while preserving inference quality and adequately propagating uncertainty across the imputed datasets.
@@
Figure 6: Number of MCMC runs required to approximate the posteriors of $m=100$ imputed datasets with $N=100$ observations, across varying model complexities ($p=10$ or $p=30$, columns) and prior specifications (standard brms priors vs. regularized horseshoe, rows). Representative datasets were selected using the Medoids procedure. The dashed horizontal reference line at 100 indicates the number of MCMC runs needed when fitting each imputed dataset separately.
Figure 6 presents the number of MCMC runs required to approximate the posteriors of all m=100{m=100} imputed datasets. The number of MCMC runs corresponds to the number of distinct proposal distributions needed for successful approximation of all target posterior distributions using our method, compared with the brute-force approach of running MCMC for all imputed datasets. The left panel of the figure shows results for the lower-dimensional case with $p=10$ predictors. Here, depending on the proportion of missing rows in the original dataset, between 1 and 55 MCMC runs are sufficient — highlighting the efficiency gains relative to running MCMC for all 100 datasets individually. As expected, the number of required MCMC runs increases with the proportion of missing data. Higher rates of missingness lead to greater variation among the imputed datasets and thus more diverse posterior distributions, making approximation more difficult for PSIS:single, PSIS:mixture and PSIS+IWMM. In the high-dimensional case with $p=30$ predictors (right panel), we observe that PSIS alone — whether based on a single proposal or a mixture — fails when the proportion of rows containing missing values is large. Nonetheless, PSIS+IWMM continues to offer substantial efficiency gains in most settings using standard priors. However, under the more challenging regularized horseshoe prior and 30% missing data, PSIS+IWMM also struggles to reduce the number of required MCMC runs effectively.
@@
Figure 7: Ratio of number of evaluations (on a log-scale) of the log-probability (left) and of the gradient log-probability (right) compared to running MCMC separately necessary to approximate the posteriors of $m=100$ imputed datasets with size $N=100$ and $p$ observations (in the columns) using either the brms standard priors or the regularized horseshoe prior (in the rows). The representative datasets are selected using the Medoids procedure.
While the previous figure focused on the number of MCMC runs, Figure 7 (left) reports the ratio of log-probability evaluations needed by the iterative algorithm relative to the number required with brute-force MCMC (baseline). A ratio close to 1, indicated by the horizontal reference line, implies that the computational effort in terms of log-probability evaluations is comparable to performing 100 separate MCMC runs. For the lower-dimensional setting (left column), the iterative algorithm requires fewer than half as many evaluations, demonstrating substantial computational savings. Results are similar for both the standard priors and the regularized horseshoe prior, though PSIS+IWMM performs slightly better with the standard priors. In contrast, for the higher-dimensional case (right column), PSIS+IWMM yields efficiency gains only when standard priors are used. In the most challenging scenario — 30% missing data and the horseshoe prior — the number of log-probability evaluations exceeds that of the brute-force MCMC baseline, indicating that the method becomes less efficient in this setting.
A similar pattern is observed when examining the ratio of gradient log-probability evaluations, as shown in Figure 7 (right). Unlike log-probability evaluations, gradient evaluations are only required during HMC sampling and are not needed for the computation of importance ratios in PSIS or IWMM. Consequently, for the inference methods based on our iterative algorithm (PSIS:single, PSIS:mixture, and PSIS+IWMM), the gradient evaluations originate exclusively from the HMC runs used to generate proposal distributions. This structural difference explains why these methods inherently require fewer or, at most, the same number of gradient evaluations as the brute-force MCMC approach. For the standard priors (top row in Figure 7 right), PSIS+IWMM shows even greater improvements in terms of gradient efficiency compared to the gains observed for log-probability evaluations. In the more challenging setting with high dimensionality and the horseshoe prior (bottom right), the performance of PSIS+IWMM is comparable to that of PSIS:single and PSIS:mixture, indicating that in this case, the added complexity of PSIS+IWMM does not lead to substantial gains.
@@
Overall, we observe that our proposed method offers substantial improvements in terms of number of log-probability and gradient evaluations — an indicator of computational cost — across most settings. For simpler priors, which result in relatively straightforward posterior distributions, the iterative approach combining both PSIS and IWMM significantly outperforms the baseline MCMC approach. Even in cases involving more complex posterior shapes, such as those induced by the regularized horseshoe prior, the proposed method continues to outperform alternative inference strategies. The only exception arises in high-dimensional settings (i.e., $p=30$ covariates) combined with a high proportion of missing data (i.e., 30% of rows with missing values), where the method does not yield improvements over the brute-force MCMC approach.
Our proposed method in the experiments with the regularized horseshoe prior performs worse than in those with the standard priors for the following reason. The stronger shrinkage induces substantially greater variation in the shape of the resulting posteriors across different datasets $D^{(i)}$. In particular, the posterior tail behavior can change markedly (e.g., heavier versus lighter tails), and in some settings the posterior can even become multimodal. As a result, posterior distributions differ much more across $D^{(i)}$, which makes approximation via PSIS and IWMM more challenging and increases the probability of failure, especially when the proposal distribution and the target posterior have fundamentally different shapes.
Figure 8: Mean absolute differences between posterior summaries obtained from MCMC and from the PSIS+IWMM resampled posterior for datasets of size $N=100$. For each missingness level (5%, 15%, and 30% of rows containing missing values), boxplots summarize the distribution across the 20 datasets and their 100 imputed datasets; within each dataset, differences are averaged over the $p$ parameters. Colors indicate the summary metric (posterior mean, standard deviation, 5% quantile, and 95% quantile). Results are shown for $p=10$ (left) and $p=30$ (right).
Although our iterative PSIS+IWMM approach reduces the number of log-probability and gradient evaluations in most settings, it approximates the posteriors via an additional step by transforming the MCMC samples with PSIS and IWMM, rather than estimating them directly with MCMC. This extra approximation layer can increase Monte Carlo standard errors and, consequently, yield slightly less accurate posterior estimates. To ensure that the approximation remains reliable, we apply the cutoff $k_{\mathrm{threshold}}=0.7$ for PSIS and IWMM diagnostics, following the recommendations of [77]. This threshold is intended to retain only cases where imprtance sampling provides a sufficiently accurate approximation to the true posterior.
To assess how closely the approximated posterior matches the MCMC baseline, we compare the resampled posteriors to the MCMC posterior using summary statistics. Figure 8 reports absolute differences in the posterior mean, standard deviation, and the 5% and 95% quantiles between PSIS+IWMM and MCMC. The reported values are averaged over the $p$ parameters. Results are shown for datasets with $N=10$, with $p=10$ parameters (left) and $p=30$ parameters (right). Overall, the posterior means are very similar, with average absolute differences below 0.0125. The same is true for the posterior standard deviations. Accurately approximating tail behavior is more challenging, as reflected by larger differences in the 5% and 95% quantiles; however, the average absolute differences remain below 0.025.
For a multivariate comparison of the posterior distributions, we additionally use the kernel maximum mean discrepancy (MMD) test [29]. This is a sample-based test with the null hypothesis that two sets of samples are drawn from the same distribution. Under a good approximation, the test should rarely reject the null at significance level $\alpha=0.05$. For example, for $N=100$, $p=10$, and 30% missingness, the null is rejected in 4.45% of cases, which is consistent with the nominal Type I error rate. In the more challenging setting with $N=10$ and a relatively large number of parameters ($p=5$), the rejection rate increases to approximately 10–15%, which is plausible given the greater estimation difficulty.
Additional results for $N=10$ and for the horseshoe prior are reported in Appendix C. We also provide results for differences between the overall posteriors obtained by pooling across the $m=100$ imputed datasets (i.e., after combining the corresponding posterior samples).
@@ 3.2 Multiple Imputation – Case Study
To evaluate the proposed method on real-world data, we examine a travel time prediction problem. Our case study is inspired by the work of [34] and utilizes a dataset provided by the NYC Taxi and Limousine Commission (TLC) [54], containing yellow taxi trip records in New York City for the year 2016. The dataset includes information such as pickup and dropoff latitude and longitude, timestamps, number of passengers, and vendor ID. Additionally, it contains variables that are only known after trip completion, such as trip distance, trip duration, and payment amount.
We focus on a short-term travel time prediction task, in line with the forecasting approach of [34]. For this purpose, we use only variables available prior to the start of the trip: pickup and dropoff latitude and longitude. In addition, we employ the Open Source Routing Machine (OSM) [19], which uses OpenStreetMap data for New York City to estimate trip duration and distance based on the trip’s start and end locations. These six variables serve as predictors in a Bayesian linear regression model, with the trip duration as the target variable. The model includes a log transformation of the target and scaling of all predictors, and we use the default priors provided by the brms [12] package.
For our analysis, we select data from a single day — Monday, March 7, 2016. The dataset is divided into 24 subsets, one for each hour of the day, based on the trip start time. Depending on the hour, the number of observations ranges from approximately 2,000 to 24,000. Each hourly subset contains some missing values, although the target variable (trip duration) is always complete. The missing data in the predictors are partly due to failures in the OSM estimation process. The proportion of missing values per hourly dataset ranges from 1.1% to 1.7%.
@@
To account for uncertainty due to missing data, we perform multiple imputation using MICE [8], generating $m=20$ imputed datasets for each hour. This choice should be sufficient, since fewer than 2% of observations are missing in the datasets. We then apply the iterative algorithm described in Section 2.3 to approximate the marginalized posterior distributions of the model parameters, using both our full approach (PSIS+IWMM) and a simplified variant without the IWMM step (PSIS:single) that employs a single proposal distribution. As demonstrated in the simulation study in Section 3.1, incorporating a mixture proposal in the multiple imputation context did not offer significant advantages over using a single proposal. Therefore, we adopted the simpler approach in this analysis. For comparison, we also perform HMC sampling (brute-force MCMC) using brms [12, 76] on each of the 20 imputed datasets individually.
Figure 9(b) presents the marginalized posterior distributions for one of these hourly datasets, specifically from 12–1 pm. The red line represents the density obtained from the brute-force MCMC approach, marginalized over all $m=20$ imputed datasets, while the blue line shows the corresponding marginalized posterior density from our iterative method using PSIS and IWMM. The close alignment in shape and location of the distributions indicates that our method closely approximates the gold-standard HMC results. Across all 24 hourly datasets and their $m=20$ imputations, the absolute difference in posterior means between the MCMC baseline and PSIS+IWMM are on the order of $10^{-4}$ when aggregated over all variables; the same holds for the posterior standard deviations. Using the MMD test [29], only 1.25% of the 480 evaluated datasets reject the null hypothesis that the MCMC and PSIS+IWMM samples are drawn from the same distribution.
@@ 3.3 Surrogate Models – Simulation Study
To evaluate the performance of the iterative algorithm in a different context, we conduct a case study based on surrogate modeling. This study replicates and extends the work of [66]. There, a two-step procedure for uncertainty quantification in surrogate models was developed and tested in multiple experiments. One of these experiments is now re-examined here using the newly developed approach.
@@ 3.3.1 Simulation Setup
$\widetilde{\mathcal{M}}_{\rm logistic}(\theta,\tau)=\frac{\tau_{1}}{1+\exp(-\tau_{2}(\theta-\tau_{3}))}+\tau_{4}\,$, (44)
@@
For the training phase of the surrogate, we assign normal priors to the surrogate parameters with means equal to the true values (2,10,0,1)(2,10,0,-1) and the standard deviations set to $(1,10,1,1)$.
@@
$\widetilde{\mathcal{M}}_{\rm PCE}(\theta,\tau)=\sum_{i=0}^{d}\tau_{i}\psi_{i}(\theta)\,$, (45)
where the parameters are $\tau=[\tau_{0},\tau_{1},\dots,\tau_{d}]$}, and $\psi_{i}$ denote the Legendre polynomials [75]. For this study, we choose a polynomial degree of $d=5$. Each coefficient $\tau_{i}$ is assigned a normal prior, $\tau_{i}\sim\mathcal{N}(0,5)$.
Surrogate Training Step In the initial step, we generate $N_{T}=10$ training data points using the simulator. The input values $\{\theta_{T}^{(n)}\}_{n=1}^{N_{T}}$ are selected as $N_{T}$ equally spaced points in the interval [1,1][-1,1]. The corresponding outputs $\{y_{T}^{(n)}\}_{n=1}^{N_{T}}$ are sampled from a normal distribution centered at the true simulator values with a standard deviation $\sigma=0.01$, i.e.,
@@
To infer the surrogate parameters, we use MCMC implemented via the probabilistic programming language Stan [76], using the cmdstanr package in R [24]. Two chains are run for 1050 iterations each (including 1000 warm-up and 50 sampling iterations), resulting in $m=100$ posterior draws $\{\tau^{(i)}\}_{i=1}^{m}\sim p(\tau\mid D_{T})$}. We use a relatively small number of posterior draws to enable a more direct comparison with the missing-data simulation study in Section 3.1. Since the choice of $m$ is difficult to determine in general, we also ran the same experiments with $m=1000$; the corresponding results are reported in Appendix D in Figure 25.
@@ 3.3.2 Results of the Simulation Study
We first examine the number of MCMC runs required by the iterative algorithm, which are used to obtain the posterior distributions serving as proposal distributions in PSIS and IWMM. Since no substantial differences between selection strategies can be observed, presented results are from the log-lik method only. Additional results for other selection methods can be found in Appendix D. Figure 10 displays the total number of MCMC runs, categorized by inference method and surrogate type (logistic and PCE). For the logistic surrogate, we observe that PSIS:single requires the highest number of MCMC runs, with a median of 37. In contrast, for PSIS:mixture only about 19 posterior distributions need to be fitted using MCMC. In most cases, the PSIS+IWMM method finishes within 1 or 2 posteriors fitted with MCMC, effectively approximating all posterior distributions and thus requiring a significantly smaller number of MCMC runs. The PCE surrogate represents a more challenging scenario because it does not include the true simulator and therefore inevitably introduces approximation error. This increases variability across the draws $\tau^{(i)}$ and, in turn, leads to greater heterogeneity in the corresponding posteriors $p(\theta\mid\tau^{(i)},D_{I})$, making approximation via PSIS and IWMM more difficult. Both methods PSIS:single and PSIS:mixture perform similarly, with a median number of MCMC runs of 70 and 63 for the single and mixture proposals, respectively. Despite some outlier cases, the combined PSIS+IWMM inference method consistently outperforms the other two in terms of minimizing the number of MCMC runs with a median of 5.
For $m=1000$ (see Figure 25 in Appendix D), the results are even more favorable. On average, fewer than 6% of the posteriors in the logistic-surrogate setting need to be computed directly via MCMC, and fewer than 20% in the PCE-surrogate setting.
Table 1: Distances between the posterior distributions $p(\theta_{I}\mid D_{T},D_{I})$ of $\theta_{I}$ obtained via MCMC and via the iterative algorithm PSIS+IWMM. Values are the absolute differences in posterior mean and standard deviation, averaged over 20 datasets.
$m$ Surrogate Mean Standard Deviation
$100$ Logistic 8.0471e-05 0.0001
PCE 0.0178 0.0528
$1000$ Logistic 0.0004 0.0005
PCE 0.0698 0.0834
@@
We also assessed how closely the resampled posterior distributions obtained with PSIS+IWMM match the gold-standard MCMC baseline. Table 1 reports the absolute differences in posterior means and standard deviations, averaged over 20 datasets. For the logistic surrogate, agreement is excellent: posterior means and standard deviations are nearly identical. In contrast, discrepancies are larger for the PCE surrogate. This is likely because the PCE surrogate has more parameters and therefore wider first-step posteriors, which propagates into greater variability in the resulting second-step posterior approximations. When comparing posterior means obtained from MCMC for the individual posteriors $p(\theta_{I}\mid\tau^{(i)},D_{I})$ across different realizations $\tau^{(i)}$, the logistic-surrogate means differ by an average absolute pairwise distance of about 0.0028 (for $m=100$). For the PCE surrogate, the corresponding average absolute difference is 0.0664. Hence, the PCE posteriors are substantially more dispersed (i.e., less similar across $\tau^{(i)}$), which helps explain why approximating them with PSIS or IWMM is more challenging.
For the logistic surrogate, the PSIS:single requires slightly more log-probability evaluations than the baseline MCMC, but significantly fewer gradient evaluations, which are typically more computationally expensive. The PSIS:mixture inference method, as well as PSIS+IWMM, succeed in approximating all posterior distributions with fewer evaluations of both log-probability and gradient evaluations. In contrast, the PCE surrogate poses a greater challenge, as the posterior distributions $p(\theta_{I}\mid\tau^{(i)},D_{I})$ differ more substantially. In the worst case — observable in PSIS:single — the iterative algorithm requires more log-probability evaluations than baseline MCMC. This occurs because the algorithm needs many iterations to evaluate all posterior distributions. During each iteration, importance ratios are computed for each of the remaining posteriors. If PSIS fails in many of these cases, a significant number of importance ratio calculations are only used to determine that PSIS does not work for this particular proposal-target combination, rather than approximating the target distribution. As a result, the number of log-probability evaluations exceeds the ones of running MCMC separately. In contrast, the number of gradient evaluations required by the new method cannot exceed those needed for running brute-force MCMC, as they are only necessary for calculating the proposal in the iterative algorithm with HMC, and there can be at most $m=100$ different proposals. The PSIS:mixture method substantially improves PSIS performance compared to the single proposal with only having a few more log-probability evaluations than the MCMC method. The combined PSIS+IWMM approach again results in the fewest gradient evaluations, indicating that only a small number of posteriors needed to be computed using HMC. The high number of log-probability evaluations in this method arises from applying the IWMM algorithm (see Section 2.2) in each iteration of the iterative algorithm for multiple different target distributions.
In conclusion, the proposed method improves the efficiency of posterior approximation for computing the marginalized posterior via the Monte Carlo estimate. In the simpler setting — using the logistic surrogate, which incorporates the true simulator — the method can approximate all posterior distributions with only one or two expensive HMC runs. In the more complex scenario involving the PCE surrogate, the method faces greater challenges, but still achieves notable reductions in computational effort. The number of gradient log-probability evaluations can be reduced to less than 14% of those required by the brute-force MCMC approach, as shown in Table 2. Since gradient evaluations typically dominate the overall computational cost, our proposed method achieves improved efficiency even in more complex settings.